OpenStudy (anonymous):

A line passes through (4, 5) and (8, 9). Which equation best represents the line? y = 5x + 8 y = 2x + 1 y = x + 1 y = 1

OpenStudy (anonymous):

Write them as functions, replace y with f(x). And substitute x with its coordinate value.

OpenStudy (anonymous):

y = f(x); solve to get the value y.

OpenStudy (anonymous):

im confused haha

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Example: y = 5x + 8 f(x) = 5x + 8 Substitute x. Since coordinates are (x, y), Either substitute x with 4 and see if the equation give out 5 or substitute x with 8 and see if the equation give out 9. Find the equation that agrees with the ordered pair.

OpenStudy (anonymous):

y = f(x)

OpenStudy (phi):

you could test if each point "works" in the choices. Only one of the choices will work. For example, the last choice y=1 means no matter what x is , y will be 1 Do your points have y=1 ?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

D?

OpenStudy (phi):

when they give you a point like (4, 5) that is a point at x=4 and y=5 it's y value is 5. 5 is not 1.

OpenStudy (anonymous):

ohhhhhh i get it now

OpenStudy (anonymous):

so it would be A

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

f(x) = y

OpenStudy (phi):

so far , you know y=1 is NOT correct (because the second number in your points is not 1)

OpenStudy (anonymous):

yea.

OpenStudy (phi):

to check choice A y = 5x + 8 use point (4,5) , that means x=4 replace x (erase it, and put in 4 ) in y = 5x + 8 then simplify can you do that ?

OpenStudy (anonymous):

13x

OpenStudy (phi):

if you replace x with 4, there won't be an "x" in the answer try again. replace x (erase it, and put in 4 ) in y = 5x + 8

OpenStudy (anonymous):

oh 5x4+8?

OpenStudy (phi):

yes, but use * for multiply (because using x is confusing) now use order of operations. do the multiply first, then add

OpenStudy (anonymous):

ok so it would be 28?

OpenStudy (phi):

yes. what you did is use the equation y= 5x+8 this equation says, "if x is 4, then you get y=28" now your point is (4,5) the point has a y=5 5 is not 28 so this is the WRONG equation. so far we crossed of A and D two more to check

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I think its B

OpenStudy (phi):

which choice do you want to check ?

OpenStudy (anonymous):

B

OpenStudy (phi):

y = 2x + 1 let's use (4,5) again. when x=4 we want to get y=5 can you test x=4 in y = 2x + 1

OpenStudy (anonymous):

=9

OpenStudy (phi):

replace x with 4 in y = 2x + 1 can you do that ?

OpenStudy (anonymous):

=9

OpenStudy (anonymous):

right? I dont know if i am doing this wrong or something?

OpenStudy (anonymous):

@phi